# When Large Language Models Meet UAVs: How Far Are We?

### v1

@@ 3.1.2. Literature Selection
Although all 997 papers retrieved in our search contained the keywords “Unmanned Aerial Vehicle (UAV)” and “Large Language Model (LLM)” in their abstracts, some of these papers did not actually discuss the integration of both technologies. For example, some studies focused on UAV systems but emphasized hardware platforms or flight control design, and were mistakenly included in the initial screening due to mentioning the technical background of LLM (Mueller, 2007). To improve the accuracy of the screening, we eliminated irrelevant articles based on the following five exclusion criteria:
@@
    Research topics that are generalized and do not focus on the combination of UAV and LLM technology;
@@
    Research that is not related to large language models;
@@
    Research that is not related to UAV technology.
To reduce bias in the filtering process, we employed a closed-card classification method to assess the relevance of the papers systematically. The classification criteria were “relevant” and “irrelevant”. The filtering was carried out independently by two graduate students, who expanded the initial dataset using snowball sampling (Bodin et al., 2019) and reviewed the abstracts of all papers. In cases of inconsistent initial judgments, the two researchers discussed and reached a consensus. After multiple rounds of filtering, we selected 74 relevant and representative papers from the initial 997 for further task categorization and quantitative analysis.
3.1.3. Data Analysis
@@
3.2. Results
Based on analysis results, we have summarized the UAV workflow into the following four steps: Information Input, Task Planning, Control Execution, and Result Feedback, which are shown in Figure 2. Each step is associated with specific tasks that LLMs assist with. In the Information Input phase, LLMs support Multimodal Information Parsing and Natural Language Command Parsing. In the Task Planning phase, LLMs are involved in Single-UAV and Multi-UAV Task Reasoning and Planning, as well as RAG-based Decision Support. The Control Execution phase utilizes Control Logic and Command Generation/Optimization. Finally, the Result Feedback phase incorporates Information Summarization and Report Generation and the development of LLM-based Chatbot.
Figure 2. Illustration of the UAV workflow and LLM task modules.
@@ 3.2. Results
Table 1. Tasks Performed by LLMs in UAV Systems.
@@
Finding 1: Academic research is distributed across the four main areas of UAV workflows: Information Input, Task Planning, Control Execution, and Result Feedback. Within these workflows, the four most common tasks are: Single-UAV Task Reasoning and Planning (39.19%), Logic and Command Generation (36.48%), Multimodal Information Parsing (31.08%), Natural Language Command Parsing (33.78%).
3.2.1. Information Input
The Information Input module is the starting point of the UAV workflow, primarily responsible for converting external commands and perception data into task information that the system can understand and execute (Colomina and Molina, 2014; Zhang et al., 2023). The tasks performed by large models in this module include the parsing of natural language commands and the analysis of multimodal data (such as images, videos, and remote sensing data), ensuring that the system can accurately understand and respond to task requirements.
Natural Language Command Parsing Natural Language Command Parsing refers to the use of large models’ natural language understanding capabilities to convert user-inputted natural language commands into structured task information. In UAV systems, before the use of LLMs, this task mainly relied on traditional speech recognition and rule-based engine technologies (Gupta et al., 2016), which often lacked sufficient understanding capabilities and could not fully analyze the context and implicit intentions of commands, leading to lower flexibility and accuracy in task execution. With LLMs, the system can accurately parse complex instructions and convert them into executable task commands through their powerful natural language understanding ability (Ahn et al., 2022). This enables UAV systems to process complex commands more intelligently, improving task execution precision and efficiency. For example, when a user inputs “Fly over the west side of the bridge and take a photo”, the system can accurately parse the flight action, target location, and task type, and generate specific execution commands. This technology has been successfully applied in systems such as Swarm-GPT (Jiao et al., 2023), TypeFly (Chen et al., 2023), and AutoHMA-LLM (Yang et al., 2025), greatly enhancing the intelligence of UAV task execution.
Multimodal Information Parsing Multimodal Information Parsing refers to the ability of large models to process non-linguistic inputs and understand scene semantics, enabling the parsing of UAV multimodal data to provide essential contextual understanding and support for task planning. In UAV systems, traditional methods relied on single sensory channels, often limiting their effectiveness in handling complex environments or multiple data sources, which led to an incomplete understanding of environmental information (Zhu et al., 2021). With the introduction of LLMs, the system can combine visual encoders (e.g., BLIP-2 (Li et al., 2023), SAM (Kirillov et al., 2023), Grounding DINO (Liu et al., 2023)) and form complete vision-language (V-L) models (Zeng et al., 2023b), allowing it to parse images, videos, and remote sensing data from UAV sensors like cameras and radar, extracting semantic information and generating language descriptions. This path significantly enhances the UAV’s ability to perceive and understand complex environments, resulting in improved performance in remote sensing, inspection, and security monitoring tasks. Representative systems such as AeroReformer (Li and Zhao, 2025), CLIPSwarm (Pueyo et al., 2024), and video anomaly detection (Zhao et al., 2024b) demonstrate the powerful capabilities of large models in multimodal data parsing, greatly enhancing UAV perception and decision-making support.
Finding 2: The highest attention in the nine tasks executed by LLMs in UAV systems is 39.19%, while the lowest is 12.16%. This difference suggests that academic attention is more evenly distributed across various tasks, with no single area of high attention.
3.2.2. Task Planning
Task Planning is primarily responsible for transforming task objectives into executable operation paths, ensuring the efficient and accurate execution of tasks. By optimizing task reasoning, path planning, and task allocation, the task planning module can handle various task requirements and complex environmental changes (Shakhatreh et al., 2019). This module involves not only the execution of single-UAV tasks but also the management of multi-UAV collaborative tasks and decision support based on external knowledge.
Single-UAV Task Reasoning and Planning Single-UAV Task Reasoning and Planning involves using large models to reason through multi-stage tasks performed by a single UAV, generating optimal paths and task execution plans. Before LLMs, this task typically relied on traditional rule-based engines or simple path-planning algorithms (Zhu et al., 2019), which struggled to adapt to dynamic changes in the environment, task priorities, and external constraints, resulting in lower task execution efficiency and flexibility. With LLMs, the system can perform real-time reasoning based on task planning, dynamically adjusting task execution plans and path planning to ensure tasks are completed in complex environments (Huang et al., 2024). This technology path enables the system to automatically adjust task order and paths based on real-time changes such as weather and obstacles. For example, when a path is affected by weather, the system will rearrange the task sequence to avoid risks. NavAgent (Liu et al., 2024) combines graph neural networks with GPT-4 for outdoor landmark-based navigation path reasoning. Additionally, reinforcement learning integration methods (Xiang et al., 2024) and the TAMP system (Sautenkov et al., 2025) further demonstrate this capability, enhancing task execution intelligence and adaptability.
Multi-UAV Task Reasoning and Planning Multi-UAV Task Reasoning and Planning refers to using large models to solve the problem of how multiple UAVs can collaborate and complete tasks. Traditional multi-UAV task scheduling relied on predefined rules and simple task allocation methods (Chung et al., 2018), which often struggled to handle dynamic environments and real-time changes, resulting in low collaboration efficiency. With LLMs, the system can intelligently allocate tasks and dynamically adjust task assignments and path planning across multiple UAVs, improving collaboration efficiency and flexibility. This technology path allows the system to adjust task areas and UAV allocations in real-time based on task priorities, environmental changes, and target status. For example, during monitoring tasks, the system automatically switches task areas based on target status, enabling efficient multi-UAV collaboration. AutoHMA-LLM (Yang et al., 2025) achieves multi-level, multi-agent collaboration, with LLM supporting dynamic task decomposition and real-time adjustments. Swarm-GPT (Jiao et al., 2023) and FlockGPT (Lykov et al., 2024) also demonstrate the effective application of this technology, significantly improving the efficiency and reliability of multi-UAV cooperative tasks.
RAG-based Decision Support RAG-based Decision Support refers to using large models combined with external knowledge to enhance the decision-making ability of UAVs. Before the use of LLMs, traditional decision support systems relied on static rule bases and limited external data, which could not flexibly handle uncertainty in complex environments. With LLMs, the system uses Retrieval-Augmented Generation (RAG) technology (Lewis et al., 2020) to quickly retrieve and analyze external data such as maps, no-fly zone information, and regulatory databases, providing precise decision support for task planning, ensuring compliance, safety, and efficiency. This technology path allows the system to assess risks in real-time during task execution and make optimal decisions. For example, SAGIN network management (Tang et al., 2024) and regulatory determination papers (Liu et al., 2024; Javaid et al., 2024) demonstrate knowledge-enhanced intelligent decision-making, improving the system’s ability to make decisions in complex environments and ensuring the successful completion of tasks.
3.2.3. Control Execution
The Control Execution module is responsible for converting the outputs of the task planning phase into specific flight control commands, ensuring that the UAV executes tasks in accordance with the requirements (Shakhatreh et al., 2019). At this stage, the system not only generates control logic and commands but also involves real-time optimization and adjustment of control strategies to cope with complex and dynamic environmental conditions.
Control Logic and Command Generation Control Logic and Command Generation refers to the use of LLMs to transform task planning results into executable flight control commands for UAVs (e.g., PX4, MAVLink), guiding UAVs to perform flight tasks. Before the introduction of LLMs, this task typically relied on manually written control logic or predefined command sets, resulting in a strong dependence on specialized programming knowledge and low efficiency in control logic generation (Koubaa et al., 2019; MAVLink Core Team, 2024). With the introduction of LLMs, the system can automatically generate flight control logic, ensuring that tasks are executed efficiently along the predetermined path. This technology path enhances the automation of control command generation, reducing human intervention and development time. For example, TypeFly (Chen et al., 2023), PX4 command generation-related papers (TAZIR et al., 2023; Zhu et al., 2024) demonstrate the automatic generation of control logic and code, effectively reducing the reliance on professional programming knowledge.
Control Logic and Command Optimization Control Logic and Command Optimization involves real-time analysis of environmental data and task feedback to optimize control strategies. Traditional methods typically rely on fixed control strategies or simple adjustment rules, which are limited in complex or dynamically changing environments (Campion et al., 2018). With LLMs, the system can assist in optimizing UAV communication strategies, link configurations, and anti-jamming designs, improving communication efficiency and stability within the swarm (Wei et al., 2025; Zhou et al., 2024). This technology path enables real-time adjustment of control logic in complex or uncertain environments, ensuring that UAVs can perform tasks efficiently and stably, especially when facing flight interference, communication issues, or other unforeseen circumstances. For example, SAGIN network management (Tang et al., 2024) demonstrates the application of LLMs in cross-layer communication network scheduling and security protection, solving resource allocation problems in complex environments. Papers (Li et al., 2025b; Yaqoot et al., 2025; Zhu et al., 2025) further provide specific implementations of this task, enhancing the effectiveness of control logic and command optimization.
3.2.4. Result Feedback
The Result Feedback module is responsible for providing information output and feedback after task execution, including task status summaries, report generation, and interaction feedback with users. This module supports real-time communication between the UAV system and the user, ensuring transparency in task execution and efficient feedback (Wang et al., 2024b; Zhang et al., 2024).
Information Summarization and Report Generation Information Summarization and Report Generation refers to the task where large models analyze UAV status, logs, perception data, and flight trajectories to generate concise and clear task reports automatically. Before the introduction of LLMs, reports were typically manually generated or relied on traditional analytical methods, which were often inefficient and lacked the flexibility to adapt to dynamic task requirements quickly. After integrating LLMs, the system can automatically generate summaries of task progress, key events, and potential risks, helping users quickly understand the task status and make timely decisions. This approach significantly improves the automation and real-time nature of task reporting. For example, after completing a monitoring task, the system can generate daily reports that include task execution status and environmental changes, enhancing the readability of task information and decision-making efficiency. Public attitude analysis (Zhu, 2024) and the Hawkeye video anomaly detection system (Zhao et al., 2024b) use LLMs to provide real-time summaries of complex data, assisting decision-makers in quickly understanding task status and risks.
LLM-based Chatbot An LLM-based Chatbot refers to the use of large models to enable natural language dialogue between humans and the UAV system. Traditional task queries and feedback were often handled through fixed interfaces or simple command responses, which lacked personalized and flexible user experiences. With the introduction of LLMs, the system can achieve task inquiries, progress feedback, and status updates through multi-turn dialogues, enhancing the intelligence and personalization of the interaction between the UAV system and the user (Zhang et al., 2024). Through this approach, users can interact more naturally with the system and quickly obtain task status updates to make real-time decisions. For example, users can ask, “How much of the current task is completed?" or "Are there any abnormalities?” The system can quickly and accurately respond, helping users stay updated on task progress. Representative UAV systems have integrated conversational interfaces for querying execution status, generating explanations, and issuing prompts (Zhao et al., 2024a; Aikins et al., 2024); in addition, mature voice and multimodal speech front-ends for UAV control can be combined with LLM dialogue further to enhance usability (Park et al., 2023; Choutri et al., 2022; Oneaţă and Cucu, 2021).
@@ 3.2.4. Result Feedback
4. RQ2: What are the differences between academia and industry in integrating UAVs with large language models?
@@
To assess the current level of industrial development in integrating LLMs with UAVs and to identify the main challenges, we compiled a dataset of real-world projects that combine these technologies. We then compared and analysed the collected papers and projects, using the four core application domains and nine task categories defined earlier, in order to clarify the gaps between academic research and industrial practice and to explore the reasons for those gaps.
4.1. Data Collection and Processing

### v2

@@ 3.1.2 Literature Selection
Although all 997 papers retrieved in our search contained the keywords “Unmanned Aerial Vehicle (UAV)” and “Large Language Model (LLM)” in their abstracts, some of these papers did not actually discuss the integration of both. For example, Li et al. [mueller2007hil]briefly introduced recent advances in LLMs in the motivation section, but the core contribution of the paper focused on designing a new UAV hardware platform and improving its flight control system. Since the study did not involve any application, evaluation, or integration of LLMs in UAV tasks, we excluded it during the final screening. To improve the accuracy of the screening, we eliminated irrelevant articles based on the following five exclusion criteria:
@@
    Research topics that are generalized and do not focus on the combination of UAVs and LLMs;
@@
    Research that is not related to LLMs;
@@
    Research that is not related to UAVs.
To reduce bias in the filtering process, we employed a closed-card classification method to assess the relevance of the papers systematically. The classification criteria were “relevant” and “irrelevant”. The filtering was carried out independently by two graduate students, who expanded the initial dataset using snowball sampling [naderifar2017snowball] and reviewed the abstracts of all papers. In cases of inconsistent initial judgments, the two researchers discussed and reached a consensus. After multiple rounds of filtering, we selected 74 relevant and representative papers from the initial 997 for further task categorization and quantitative analysis.
3.1.3 Data Analysis
@@
3.2 Results
Based on analysis results, we have summarized the UAV workflow into the following four steps: Information Input, Task Planning, Control Execution, and Result Feedback, which are shown in Figure LABEL:fig:fig4. Each step is associated with specific tasks that LLMs assist with. In the Information Input phase, LLMs support Multimodal Information Parsing and Natural Language Command Parsing. In the Task Planning phase, LLMs are involved in Single-UAV and Multi-UAV Task Reasoning and Planning, as well as RAG-based Decision Support. The Control Execution phase utilizes Control Logic and Command Generation/Optimization. Finally, the Result Feedback phase incorporates Information Summarization and Report Generation and the development of LLM-based Chatbot.
Figure 2: Illustration of the UAV workflow and LLM task modules.
@@ 3.2 Results
Table 1: Tasks Performed by LLMs in UAV Systems.
@@
[…]
3.2.1 Information Input
The Information Input module is the starting point of the UAV workflow, primarily responsible for converting external commands and perception data into task information that the system can understand and execute [colomina2014unmanned, survey2023uavedge]. The tasks performed by large models in this module include the parsing of natural language commands and the analysis of multimodal data (such as images, videos, and remote sensing data), ensuring that the system can accurately understand and respond to task requirements.
Natural Language Command Parsing. Natural Language Command Parsing refers to the use of large models’ natural language understanding capabilities to convert user-inputted natural language commands into structured task information. In UAV systems, before the use of LLMs, this task mainly relied on traditional speech recognition and rule-based engines [gupta2016survey], which often lacked sufficient understanding capabilities and could not fully analyze the context and implicit intentions of commands, leading to lower flexibility and accuracy in task execution. With LLMs, the system can accurately parse complex instructions and convert them into executable task commands through their powerful natural language understanding ability [ahn2022saycan]. This enables UAV systems to process complex commands more intelligently, improving task execution precision and efficiency. For example, when a user inputs “Fly over the west side of the bridge and take a photo”, the system can accurately parse the flight action, target location, and task type, and generate specific execution commands. This has been successfully applied in systems such as Swarm-GPT [jiao2023swarm], TypeFly [chen2023typefly], and AutoHMA-LLM [10839354], greatly enhancing the intelligence of UAV task execution.
Multimodal Information Parsing. Multimodal Information Parsing refers to the ability of large models to process non-linguistic inputs and understand scene semantics, enabling the parsing of UAV multimodal data to provide essential contextual understanding and support for task planning. In UAV systems, traditional methods relied on single sensory channels, often limiting their effectiveness in handling complex environments or multiple data sources, which led to an incomplete understanding of environmental information [zhu2021survey]. With the introduction of LLMs, the system can combine visual encoders (e.g., BLIP-2 [li2023blip2], SAM [kirillov2023sam], Grounding DINO [liu2023groundingdino]) and form complete vision-language (V-L) models [zeng2023vlmsurvey], allowing it to parse images, videos, and remote sensing data from UAV sensors like cameras and radar, extracting semantic information and generating language descriptions. This path significantly enhances the UAV’s ability to perceive and understand complex environments, resulting in improved performance in remote sensing, inspection, and security monitoring tasks. Representative systems such as AeroReformer [li2025aeroreformeraerialreferringtransformer], CLIPSwarm [10801327], and video anomaly detection [zhao2024hawkeye] demonstrate the powerful capabilities of large models in multimodal data parsing, greatly enhancing UAV perception and decision-making support.
3.2.2 Task Planning
Task Planning is primarily responsible for transforming task objectives into executable operation paths, ensuring the efficient and accurate execution of tasks. By optimizing task reasoning, path planning, and task allocation, the task planning module can handle various task requirements and complex environmental changes [shakhatreh2019unmanned]. This module involves not only the execution of single-UAV tasks but also the management of multi-UAV collaborative tasks and decision support based on external knowledge.
Single-UAV Task Reasoning and Planning. Single-UAV Task Reasoning and Planning involves using large models to reason through multi-stage tasks performed by a single UAV, generating optimal paths and task execution plans. Before LLMs, this task typically relied on traditional rule-based engines or simple path-planning algorithms [zhu2019uavpathplanning], which struggled to adapt to dynamic changes in the environment, task priorities, and external constraints, resulting in lower task execution efficiency and flexibility. With LLMs, the system can perform real-time reasoning based on task planning, dynamically adjusting task execution plans and path planning to ensure tasks are completed in complex environments [huang2024lmplanners]. This approach enables the system to automatically adjust task order and paths based on real-time changes such as weather and obstacles. For example, when a path is affected by weather, the system will rearrange the task sequence to avoid risks. NavAgent [liu2024navagent] combines graph neural networks with GPT-4 for outdoor landmark-based navigation path reasoning. Additionally, reinforcement learning integration methods [10650538] and the TAMP system [sautenkov2025uavvlavisionlanguageactionlargescale] further demonstrate this capability, enhancing task execution intelligence and adaptability.
Multi-UAV Task Reasoning and Planning. Multi-UAV Task Reasoning and Planning refers to using large models to solve the problem of how multiple UAVs can collaborate and complete tasks. Traditional multi-UAV task scheduling relied on predefined rules and simple task allocation methods [chung2018multiuavsurvey], which often struggled to handle dynamic environments and real-time changes, resulting in low collaboration efficiency. With LLMs, the system can intelligently allocate tasks and dynamically adjust task assignments and path planning across multiple UAVs, improving collaboration efficiency and flexibility. This approach allows the system to adjust task areas and UAV allocations in real-time based on task priorities, environmental changes, and target status. For example, during monitoring tasks, the system automatically switches task areas based on target status, enabling efficient multi-UAV collaboration. AutoHMA-LLM [10839354] achieves multi-level, multi-agent collaboration, with LLM supporting dynamic task decomposition and real-time adjustments. Swarm-GPT [jiao2023swarm] and FlockGPT [10765199] significantly improve the efficiency and reliability of multi-UAV cooperative tasks.
RAG-based Decision Support. RAG-based Decision Support refers to using LLMs combined with external knowledge to enhance the decision-making ability of UAVs. Before the use of LLMs, traditional decision support systems relied on static rule bases and limited external data, limiting their flexibility in uncertain environments. With LLMs, the system utilizes Retrieval-Augmented Generation (RAG) [lewis2020rag] to retrieve and analyze vast amounts of external data, including maps, no-fly zone information, and regulatory databases. It can inform the planning process, ensuring that the plans generated comply with safety standards and the latest intelligence. For example, in SAGIN network management [69] and regulatory compliance scenarios [liu2024navagent, javaid2024large], RAG-based support enables the system to incorporate domain-specific knowledge into decision-making, ensuring that complex tasks are not only logically sound but also contextually accurate and compliant.
3.2.3 Control Execution
The Control Execution module is responsible for converting the outputs of the task planning phase into specific flight control commands, ensuring that the UAV executes tasks in accordance with the requirements [shakhatreh2019unmanned]. At this stage, the system not only generates control logic and commands but also involves real-time optimization and adjustment of control strategies to cope with complex and dynamic environmental conditions.
Control Logic and Command Generation. Control Logic and Command Generation refers to the use of LLMs to transform task planning results into executable flight control commands for UAVs (e.g., PX4, MAVLink), guiding UAVs to perform flight tasks. Before the introduction of LLMs, this task typically relied on manually written control logic or predefined command sets, resulting in a strong dependence on specialized programming knowledge and low efficiency in control logic generation [koubaa2019mavlink, mavlinkdevguide]. With the introduction of LLMs, the system can automatically generate flight control logic, ensuring that tasks are executed efficiently along the predetermined path. This approach enhances the automation of control command generation, reducing human intervention and development time. For example, TypeFly [chen2023typefly], PX4 command generation-related papers [tazir2023words, 56] demonstrate the automatic generation of control logic and code, effectively reducing the reliance on professional programming knowledge.
Control Logic and Command Optimization. Control Logic and Command Optimization involves real-time analysis of environmental data and task feedback to optimize control strategies. Traditional methods typically rely on fixed control strategies or simple adjustment rules, which are limited in complex or dynamically changing environments [campion2018swarmsurvey]. With LLMs, the system can assist in optimizing UAV communication strategies, link configurations, and anti-jamming designs, improving communication efficiency and stability within the swarm [wei2025llmwirelesssurvey, zhou2024llmicl]. This approach enables real-time adjustment of control logic in complex or uncertain environments, ensuring that UAVs can perform tasks efficiently and stably, especially when facing flight interference, communication issues, or other unforeseen circumstances. For example, SAGIN network management [69] demonstrates the application of LLMs in cross-layer communication network scheduling and security protection, solving resource allocation problems in complex environments. Papers [15, 30, 50] further provide specific implementations of this task, enhancing the effectiveness of control logic and command optimization.
3.2.4 Result Feedback
The Result Feedback module is responsible for providing information output and feedback after task execution, including task status summaries, report generation, and interaction feedback with users. This module supports real-time communication between the UAV system and the user, ensuring transparency in task execution and efficient feedback [wang2024mhri, zhang2024llmhri].
Information Summarization and Report Generation. Information Summarization and Report Generation refers to the task where large models analyze UAV status, logs, perception data, and flight trajectories to generate concise and clear task reports automatically. Before the introduction of LLMs, reports were typically manually generated or relied on traditional analytical methods, which were often inefficient and lacked the flexibility to adapt to dynamic task requirements quickly. After integrating LLMs, the system can automatically generate summaries of task progress, key events, and potential risks, helping users quickly understand the task status and make timely decisions. This approach significantly improves the automation and real-time nature of task reporting. For example, after completing a monitoring task, the system can generate daily reports that include task execution status and environmental changes, enhancing the readability of task information and decision-making efficiency. Public attitude analysis [23] and the Hawkeye video anomaly detection system [zhao2024hawkeye] use LLMs to provide real-time summaries of complex data, assisting decision-makers in quickly understanding task status and risks.
LLM-based Chatbot. LLM-based Chatbot refers to the use of large models to enable natural language dialogue between humans and the UAV system. Traditional task queries and feedback were often handled through fixed interfaces or simple command responses, which lacked personalized and flexible user experiences. With the introduction of LLMs, the system can achieve task inquiries, progress feedback, and status updates through multi-turn dialogues, enhancing the intelligence and personalization of the interaction between the UAV system and the user [zhang2024llmhri]. Through this approach, users can interact more naturally with the system and quickly obtain task status updates to make real-time decisions. For example, users can ask, “How much of the current task is completed?” or ”Are there any abnormalities?” The system can quickly and accurately respond, helping users stay updated on task progress. Representative UAV systems have integrated conversational interfaces for querying execution status, generating explanations, and issuing prompts [zhao2024multibotgpt, aikins2024leviosa]; in addition, mature voice and multimodal speech front-ends for UAV control can be combined with LLM dialogue further to enhance usability [park2023voiceuav, choutri2022multilingual, oneata2021multimodal].
@@ 3.2.4 Result Feedback
4 RQ2: What are the differences between researchers’ and developers’ practices in integrating UAV projects with large language models?
@@
To assess the current level of developers’ practices in integrating LLMs with UAVs and to identify the main challenges, we compiled a dataset of real-world projects that combine them. We then compared and analysed the collected papers and projects, using the four core application domains and nine task categories defined earlier, in order to clarify the gaps between researchers’ and developers’ practices and to explore the reasons for those gaps.
4.1 Data Collection and Processing